# There is no universal separable Banach algebra

### v1

@@ Abstract.
We show that no separable Banach algebra is universal for homomorphic embeddings of all separable Banach algebras, whether embeddings are merely bounded or required to be contractive. The commutative version also fails: no separable commutative Banach algebra is universal for embeddings of all separable commutative algebras.
The proofs follow the same pattern but use different linearisation spaces. Given a bounded bilinear form $\beta$, we associate a separable test algebra $A(\beta)$ whose multiplication records $\beta$. Any homomorphic embedding of $A(\beta)$ forces the linearisation of $\beta$ to factor through a fixed separable space: in the commutative case through $R\widehat{\otimes}_{\pi,s}R$ (where $R$ is the Jacobson radical of the target algebra), and in the general case through $B\widehat{\otimes}_{\pi}B$. Choosing $\beta$ so that the corresponding operator fails to factor through that space, using the theorem of Johnson–Szankowski, yields a contradiction. The non‑commutative argument avoids radicals entirely and uses only the ordinary projective tensor product.
@@
Banach algebra, Jacobson radical, symmetric projective tensor product, universal Banach algebra, Johnson–Szankowski theorem
@@ 1. Introduction
For commutative C*-algebras, the situation is well understood. Since every compact metrisable space is a quotient of the Cantor set $\Delta:=2^{\mathbb{N}}$, the algebra $C(\Delta)$ of continuous functions on the Cantor set is universal for separable commutative C*-algebras: every such algebra embeds isometrically as a subalgebra of $C(\Delta)$. In contrast, there is no separable C*-algebra that is universal for all separable C*-algebras (commutative or not), as can be deduced from [5, Proposition 2.6].
@@
Our main result, Theorem 1.1, shows that there is no separable (commutative) Banach algebra $B$ that is universal for embeddings of separable (commutative) Banach algebras, both in the category of contractive homomorphisms and in the category of (arbitrary) bounded homomorphisms. The proofs combine constructions using projective tensor products (symmetric in the commutative case, ordinary in the general case) with the deep Johnson–Szankowski theorem [4, Theorem B] that the class of compact operators does not factor through any separable Banach space. By constructing test algebras whose radical structure forces any homomorphic embedding to induce a factorisation of a carefully chosen non-factorable compact operator, we obtain the desired contradiction.
@@
We present a single proof that treats the commutative and general cases separately but in either case covers both morphism kinds at once. Contractivity plays no rôle beyond boundedness.
@@ 2. Preliminaries
We refer to [7, Chapter 2] for all facts related to the projective tensor product of Banach spaces and to [1, Chapter 2] for Banach-algebraic aspects thereof. For a Banach space $X$ and an integer $n$, the ($n$-fold) projective tensor product $X^{\widehat{\otimes}_{\pi}n}$ (or simply $X\widehat{\otimes}_{\pi}X$ for $n=2$) is defined as the completion of the algebraic tensor product $X^{\otimes n}$ under the projective norm.
@@
The well-defininedness of the operator $u_{1}\otimes_{\pi}u_{2}$ is standard consequence of the universality of the projective tensor product and the norm estimate follows from reasonability of the projective crossnorm. For the identity, evaluate both sides on elementary tensors:
@@
In particular, if $M=N$ as bilinear maps on the same domain, then $\widetilde{M}=\widetilde{N}$.
@@
Let $u:X\to Z$ be bounded. Then there is a unique bounded operator
$u\otimes_{\mathrm{sym}}u:\ X\widehat{\otimes}_{\pi,s}X\ \longrightarrow\ Z\widehat{\otimes}_{\pi,s}Z$
with $(u\otimes_{\mathrm{sym}}u)(x\odot x^{\prime})=u(x)\odot u(x^{\prime})$ on symmetric tensors, and
$\|\,u\otimes_{\mathrm{sym}}u\,\|\ \leqslant\ \|u\|^{2}.$
@@
Let $q_{X}:X\widehat{\otimes}_{\pi}X\to X\widehat{\otimes}_{\pi,s}X$ and $q_{Z}:Z\widehat{\otimes}_{\pi}Z\to Z\widehat{\otimes}_{\pi,s}Z$ be the canonical quotient maps (of norm $1$). Set $T:=q_{Z}\circ(u\otimes_{\pi}u)$; then $\|T\|\leqslant\|u\|^{2}$ by Lemma 2.2. By the universal property of $q_{X}$ there exists a unique $U:X\widehat{\otimes}_{\pi,s}X\to Z\widehat{\otimes}_{\pi,s}Z$ with $U\circ q_{X}=T$; this is $u\otimes_{\mathrm{sym}}u$. For $z\in X\widehat{\otimes}_{\pi,s}X$ and $\varepsilon>0$ choose $w$ with $q_{X}(w)=z$ and $\|w\|\leqslant\|z\|+\varepsilon$. Then
$\|Uz\|=\|Tw\|\leqslant\|T\|\,\|w\|\leqslant\|u\|^{2}(\|z\|+\varepsilon);$
letting $\varepsilon\downarrow 0$ proves the claim. ∎
For a Banach space $X$ and an integer $n$, the ($n$-fold) symmetric projective tensor product $X^{\widehat{\otimes}_{\pi,s}n}$ (or simply $X\widehat{\otimes}_{\!\pi,s}X$ for $n=2$) is defined as the completion of the algebraic symmetric tensor product $X^{\odot n}$ under the projective norm inherited from $X^{\widehat{\otimes}_{\pi}n}$. Here $X^{\odot n}$ denotes the quotient of the algebraic tensor product $X^{\otimes n}$ by the subspace generated by all elements of the form $x_{1}\otimes\cdots\otimes x_{n}-x_{\sigma(1)}\otimes\cdots\otimes x_{\sigma(n)}$, with $\sigma$ ranging over the symmetric group $S_{n}$. It provides a natural linearisation space for continuous symmetric $n$-linear maps.
Theorem 2.5 (Universality of the symmetric projective tensor product).
Let $X,Y$ be Banach spaces and $n\in\mathbb{N}$. For every continuous symmetric $n$-linear map $M\colon X^{n}\to Y$ there exists a unique bounded operator $L_{M}\in\mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi,s}n},\,Y\big)$ such that
$M=L_{M}\circ\otimes^{n}_{s},\qquad\text{equivalently}\quad M=(L_{M}\circ\sigma^{n}_{X})\circ\otimes^{n}_{\pi}.$
Here $\otimes^{n}_{\pi}\colon X^{n}\to X^{\widehat{\otimes}_{\pi}n}$ is the canonical $n$-linear map, $\sigma^{n}_{X}\colon X^{\widehat{\otimes}_{\pi}n}\to X^{\widehat{\otimes}_{\pi,s}n}$ is the canonical quotient onto the symmetric projective tensor product, and $\otimes^{n}_{s}:=\sigma^{n}_{X}\circ\otimes^{n}_{\pi}$. Diagrammatically,
[…]
$\mathcal{L}_{s}(X^{n};Y)\ \cong\ \mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi,s}n},\,Y\big),$
so that $\|L_{M}\|=\|M\|$. In particular, $(X^{\widehat{\otimes}_{\pi,s}n})^{*}$ is isometrically the space of continuous $n$-homogeneous polynomials on $X$.
Proof.
This is standard; see Hájek–Johanis [2, Theorem 13]. ∎
Remark 2.6.
For any bounded bilinear $\beta:X\times X\to Y$ we write $\widetilde{\beta}:X\widehat{\otimes}_{\pi}X\to Y$ for its (unsymmetrised) projective linearisation. If $\beta$ is symmetric, we also write $\widetilde{\beta}^{\,\mathrm{sym}}:X\widehat{\otimes}_{\pi,s}X\to Y$ for the symmetric linearisation. In the commutative parts of the paper we use $\widetilde{\beta}^{\,\mathrm{sym}}$ consistently.
If $R$ is a commutative Banach algebra, its multiplication $\mu_{R}:R\times R\to R$ is symmetric and linearises to $\widetilde{\mu}_{R}^{\mathrm{sym}}:R\widehat{\otimes}_{\!\pi,s}R\to R$.
Lemma 2.7 (Image of the radical under homomorphisms).
Let $\phi:A\to B$ be a bounded homomorphism of Banach algebras. Then $\phi\bigl(\operatorname{rad}(A)\bigr)\subseteq\operatorname{rad}\bigl(\overline{\phi(A)}\bigr)$.
Proof.
This is standard: the Jacobson radical is functorial for surjective maps and, in general, one has $\phi(\operatorname{rad}A)\subseteq\operatorname{rad}(\overline{\phi(A)})$; see, e.g., [1, §3.3]. ∎
Lemma 2.8.
Let $A$ be a Banach algebra and let $I\triangleleft A$ be a (two-sided) ideal with $I^{n}=0$ for some $n\in\mathbb{N}$. Then $I\subseteq\operatorname{rad}(A)$. Moreover, if $A/I$ is semisimple (equivalently $\operatorname{rad}(A/I)=0$), then $\operatorname{rad}(A)\subseteq I$.
Proof.
If $a\in I$ then $(1-a)(1+a+\dots+a^{n-1})=1$, so 1a1-a is invertible in $A$. Hence $a$ lies in the Jacobson radical. The second assertion follows from the standard functoriality of the radical under quotient maps, see [1, §3.3]. ∎
We now construct our test algebras. Given a bounded bilinear map $\beta:X\times X\to Y$, we build a three-level algebra $A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y$ where the scalar level $\mathbb{C}$ acts as an identity, the middle level $X$ acts on itself via $\beta$ to produce elements in the top level $Y$, and the top level $Y$ is multiplicatively inert (squares to zero). The multiplication thus ‘records’ the bilinear form $\beta$. Any homomorphism embedding $A(\beta)$ into another algebra will be forced to preserve this structure, leading to factorisation constraints that we exploit via the Johnson–Szankowski theorem.
@@
If $\beta$ is symmetric, it linearises to a unique bounded operator $\widetilde{\beta}:X\widehat{\otimes}_{\pi,\mathrm{s}}X\to Y$.
@@
This argument uses only that $J$ is nilpotent, so it holds whether or not $\beta$ is symmetric.
Finally, we observe that the symmetric bilinear map $\beta:X\times X\to Y$ linearises, by Theorem 2.5, to a unique bounded linear operator $\widetilde{\beta}:X\widehat{\otimes}_{\pi,\mathrm{s}}X\to Y$ satisfying $\widetilde{\beta}(x\otimes_{\mathrm{sym}}x^{\prime})=\beta(x,x^{\prime})$ for all $x,x^{\prime}\in X$, with $\|\widetilde{\beta}\|=\|\beta\|$. ∎
We now show that when $A(\beta)$ embeds into a commutative Banach algebra $B$, the image must respect the radical structure. Since $\operatorname{rad}(A(\beta))=X\oplus Y$ is nilpotent and homomorphisms preserve radicals in the commutative setting, the embedding forces $X$ and $Y$ to land in $\operatorname{rad}(B)$. The multiplication structure in $A(\beta)$ then translates into a factorisation identity through the symmetric tensor product of the radical.
Lemma 2.10.
Let $B$ be a commutative Banach algebra with radical $R=\operatorname{rad}(B)$, and let $X$ and $Y$ be Banach spaces. Consider the algebra
$A(\beta):=\mathbb{C}\oplus_{1}X\oplus_{1}Y,\qquad(\alpha,x,y)\cdot(\alpha^{\prime},x^{\prime},y^{\prime})=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr),$
where $\beta:X\times X\to Y$ is a bounded symmetric bilinear map, and let $\widetilde{\beta}^{\,\mathrm{sym}}:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to Y$ be its unique bounded linearisation through the symmetric projective tensor product. Suppose that $\phi:A(\beta)\to B$ is a bounded algebra homomorphism, and write
$i:=\phi|_{X}:X\longrightarrow R,\qquad j:=\phi|_{Y}:Y\longrightarrow R.$
Then
(2.1) $j\circ\widetilde{\beta}^{\,\mathrm{sym}}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(\,i\otimes_{\mathrm{sym}}i\,)\quad:\quad X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\longrightarrow R,$
where $\mu_{R}:R\times R\to R$ denotes multiplication in the radical $R$, $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R\to R$ is its symmetric linearisation, and $i\otimes_{\mathrm{sym}}i:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R$ is the bounded map induced by $i$ on symmetric tensor products. Moreover, if $\phi$ is a homomorphic embedding, then both $i$ and $j$ are injective, and $j$ is bounded below. Hence $j$ (and similarly $i$) is injective with closed range and admits a bounded inverse on its range.
Proof.
By Proposition 2.9, $J:=X\oplus Y$ is an ideal of $A(\beta)$ with $J^{3}=0$, hence every $z\in J$ is nilpotent. Therefore $\phi(z)$ is nilpotent in $B$ and lies in $R:=\operatorname{rad}(B)$. In particular the restrictions
$i:=\phi|_{X}:X\to R,\qquad j:=\phi|_{Y}:Y\to R$
are well-defined and bounded.
For each $x\in X$, write $\mathbf{x}:=(0,x,0)\in A(\beta)$, and for each $y\in Y$ write $\mathbf{y}:=(0,0,y)\in A(\beta)$. The multiplication in $A(\beta)$ gives
$\mathbf{x}\cdot\mathbf{x}^{\prime}\ =\ (0,0,\beta(x,x^{\prime}))\qquad(x,x^{\prime}\in X).$
Applying the homomorphism $\phi$ to this identity, and noting that $\phi(\mathbf{x})=i(x)$, $\phi(\mathbf{x}^{\prime})=i(x^{\prime})$, and $\phi(0,0,y)=j(y)$ all lie in $R$, we obtain
(2.2) $j\bigl(\beta(x,x^{\prime})\bigr)\ =\ \phi(\mathbf{x}\cdot\mathbf{x}^{\prime})\ =\ \phi(\mathbf{x})\,\phi(\mathbf{x}^{\prime})\ =\ \mu_{R}\bigl(i(x),i(x^{\prime})\bigr)$
for all $x,x^{\prime}\in X$.
As $\beta$ is a bounded symmetric bilinear map, we may apply Theorem 2.5 with $Z=Y$ to obtain a unique bounded operator $\widetilde{\beta}:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to Y$ satisfying $\widetilde{\beta}(x\odot_{\mathrm{s}}x^{\prime})=\beta(x,x^{\prime})$ for all $x,x^{\prime}\in X$. Similarly, since $R$ is commutative, the multiplication $\mu_{R}:R\times R\to R$ is a bounded symmetric bilinear map. Applying Theorem 2.5 with $Z=R$, we obtain a unique bounded operator $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R\to R$ such that $\widetilde{\mu}_{R}^{\,\mathrm{sym}}(u\odot_{\mathrm{s}}v)=\mu_{R}(u,v)$ for all $u,v\in R$.
Furthermore, by Lemma 2.7, the homomorphism $\phi:A(\beta)\to B$ maps $\operatorname{rad}(A(\beta))$ into the radical $R=\operatorname{rad}(B)$, so its restriction $i=\phi|_{X}:X\to R$ is a bounded linear map. The functoriality of the symmetric projective tensor product then yields a canonical bounded operator
$i\otimes_{\mathrm{sym}}i\colon X\widehat{\otimes}_{\pi,\mathrm{s}}X\longrightarrow R\widehat{\otimes}_{\pi,\mathrm{s}}R,$
defined on elementary symmetric tensors by
$(i\otimes_{\mathrm{sym}}i)(x\odot_{\mathrm{s}}x^{\prime})=i(x)\odot_{\mathrm{s}}i(x^{\prime}),\qquad x,x^{\prime}\in X,$
and extended to the completion by continuity. The map $i\mapsto i\otimes_{\mathrm{sym}}i$ is continuous and one has the norm estimate $\|\,i\otimes_{\mathrm{sym}}i\,\|\leqslant\|i\|^{2}$.
Consider the bounded operators
$T_{1}:=j\circ\widetilde{\beta}:\ X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\ \to\ R,\qquad T_{2}:=\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i):\ X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\ \to\ R.$
For all $x,x^{\prime}\in X$, we have by construction and by (2.2):
$T_{1}(x\odot_{\mathrm{s}}x^{\prime})\$$=\ j\bigl(\beta(x,x^{\prime})\bigr)\ =\ \mu_{R}\bigl(i(x),i(x^{\prime})\bigr)$
$=\ \widetilde{\mu}_{R}^{\,\mathrm{sym}}\bigl(i(x)\odot_{\mathrm{s}}i(x^{\prime})\bigr)\ =\ T_{2}(x\odot_{\mathrm{s}}x^{\prime}).$
[…]
Suppose now that $\phi$ is an embedding, so that $\phi$ is injective and admits a bounded inverse $\psi:\phi(A(\beta))\to A(\beta)$ on its range. Then, for each $y\in Y$, we have
$\|y\|\ =\ \bigl\|\psi\bigl(\phi(0,0,y)\bigr)\bigr\|\ \leqslant\ \|\psi\|\,\bigl\|\phi(0,0,y)\bigr\|\ =\ \|\psi\|\,\|j(y)\|.$
Consequently, $\|j(y)\|\geqslant\|\psi\|^{-1}\|y\|$ for all $y\in Y$, which shows that $j$ is bounded below. In particular, $j$ is injective. An entirely analogous argument applied to $x\in X$ establishes that $i=\phi|_{X}$ is injective (and likewise bounded below). ∎
The proof works the same for the unsymmetrised version.
@@
Let $M:=\mu_{B}\circ(i\times i)$ and $N:=j\circ\beta$ as bounded bilinear maps on $X\times X$. From $(0,x,0)(0,x^{\prime},0)=(0,0,\beta(x,x^{\prime}))$ in $A(\beta)$ and multiplicativity of $\phi$ we get $M(x,x^{\prime})=N(x,x^{\prime})$ for all $x,x^{\prime}$. By Lemma 2.3 we obtain $\widetilde{M}=\widetilde{N}$. Using Lemma 2.2, $\widetilde{M}=\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i)$, while $\widetilde{N}=j\circ\widetilde{\beta}$, yielding the identity. If $\phi$ is an embedding with inverse $\psi$ on its range, then $\|y\|=\|\psi(\phi(0,0,y))\|\leqslant\|\psi\|\,\|j(y)\|$, so $j$ is bounded below. ∎
Remark 2.12.
Note that Lemma 2.11 does not use any radical structure: $i$ and $j$ need not land in a distinguished ideal of $B$. All factorisations proceed through the fixed separable space
$G:=B\widehat{\otimes}_{\pi}B$
via the canonical map $i\otimes_{\pi}i:E\to G$ and the linearised multiplication $\widetilde{\mu}_{B}:G\to B$.
To apply the Johnson–Szankowski theorem, we must carefully choose the bilinear form $\beta$ so that its linearisation $\widetilde{\beta}$ does not factor through a given separable space $G$. The construction proceeds in two steps: first, we build a surjective bilinear form $\beta_{U}$ with a bounded right inverse (at the linearised level), and then compose with a compact operator $S:U\to Y$ that fails to factor through $G$. The surjectivity ensures that non-factorisation of the composition implies non-factorisation of $S$, completing the circle.
@@
Let $G$ be a separable Banach space, and let $U$ and $Y$ be Banach spaces. Suppose that $S:U\to Y$ is a compact operator which does not factor through $G$. Write $X:=U\oplus_{1}\mathbb{C}$, and define a symmetric bilinear map $\beta_{U}:X\times X\to U$ by
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big)=\lambda u^{\prime}+\lambda^{\prime}u$
for $(u,\lambda),(u^{\prime},\lambda^{\prime})\in X$. Let $E:=X\widehat{\otimes}_{\pi,\mathrm{s}}X$, and let $\widetilde{\beta}_{U}:E\to U$ be the linearisation of $\beta_{U}$. Then $\widetilde{\beta}_{U}$ is surjective and admits a bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$. Consequently, writing $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, we have:
  • $T_{0}$ is compact, and
  • $T_{0}$ does not factor through $G$ (for otherwise $S=T_{0}\circ\sigma$ would factor through $G$, contradicting our hypothesis).
Proof.
For each $u\in U$, observe that
$\widetilde{\beta}_{U}\big((u,0)\otimes_{\mathrm{sym}}(0,1)\big)=\beta_{U}\big((u,0),(0,1)\big)=u.$
Thus $\widetilde{\beta}_{U}$ is surjective, and the map $\sigma:U\to E$ defined by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$ satisfies $\widetilde{\beta}_{U}\circ\sigma=\operatorname{id}_{U}$.
To verify that $\sigma$ is bounded, for elementary symmetric tensors in any symmetric projective tensor product, one has
$\big\|x\otimes_{\mathrm{sym}}y\big\|_{\pi,\mathrm{s}}\leqslant\|x\|\,\|y\|,$
because $X\widehat{\otimes}_{\!\pi,s}X$ is a quotient of the projective tensor product that is a reasonable crossnorm. Consequently,
$\|\sigma(u)\|_{\pi,\mathrm{s}}\ \leqslant\ \|(u,0)\|\,\|(0,1)\|\ =\ \|u\|\cdot 1\ =\ \|u\|,$
so that $\sigma$ is indeed bounded (in fact, a contraction).
Since $S$ is compact and $\widetilde{\beta}_{U}$ is bounded, their composition $T_{0}=S\circ\widetilde{\beta}_{U}$ is compact. Suppose, towards a contradiction, that $T_{0}$ factors through $G$, say $T_{0}=Q\circ P$ with $P:E\to G$ and $Q:G\to Y$ both bounded. Then
$S\ =\ T_{0}\circ\sigma\ =\ Q\circ(P\circ\sigma),$
which exhibits $S$ as factoring through $G$, contrary to hypothesis. Thus $T_{0}$ does not factor through $G$. ∎
An analogous fact holds true in the non-symmetric setting.
Lemma 2.14.
Let $G$ be a separable Banach space, and let $U,Y$ be Banach spaces. Suppose $S:U\to Y$ is a compact operator which does not factor through $G$. Set $X:=U\oplus_{1}\mathbb{C}$ and define a bilinear map
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda\,u^{\prime}\qquad\big((u,\lambda),(u^{\prime},\lambda^{\prime})\in X\big).$
Let $E:=X\widehat{\otimes}_{\pi}X$, and let $\widetilde{\beta}_{U}:E\to U$ be its linearisation. Then $\widetilde{\beta}_{U}$ is surjective and admits a bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(0,1)\otimes(u,0)$. Consequently, writing $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, we have:
@@
For each $u\in U$,
$\widetilde{\beta}_{U}\big((0,1)\otimes(u,0)\big)=\beta_{U}\big((0,1),(u,0)\big)=u,$
so $\widetilde{\beta}_{U}$ is surjective and $\sigma(u):=(0,1)\otimes(u,0)$ is a right inverse. Since $\|(0,1)\otimes(u,0)\|_{\pi}\leqslant\|(0,1)\|\,\|(u,0)\|=\|u\|$, the map $\sigma$ is bounded. Compactness of $T_{0}=S\circ\widetilde{\beta}_{U}$ follows because $S$ is compact and $\widetilde{\beta}_{U}$ is bounded. If $T_{0}$ factored through $G$ as $Q\circ P$ then $S=T_{0}\circ\sigma=Q\circ(P\circ\sigma)$ would factor through $G$, contradicting the hypothesis. ∎

### v2

@@ Abstract.
We prove that no separable Banach algebra is universal for homomorphic embeddings of all separable Banach algebras, whether embeddings are merely bounded or required to be contractive. The same holds in the commutative category.
The proof uses the following scheme. To each bounded bilinear form $\beta$ we attach a separable test algebra $A(\beta)$ whose multiplication records $\beta$. Any homomorphic embedding of $A(\beta)$ into a candidate $B$ forces the linearisation of $\beta$ to factor through the fixed separable space $B\widehat{\otimes}_{\pi}B$. Choosing $\beta$ so that the associated operator fails to factor through $B\widehat{\otimes}_{\pi}B$, by the theorem of Johnson–Szankowski, yields a contradiction. In the commutative case, we take $\beta$ symmetric so $A(\beta)$ is commutative.
@@
Banach algebra, projective tensor product, factorisation of operators, universal Banach algebra, Johnson–Szankowski theorem
@@ 1. Introduction
For commutative C*-algebras, the situation is well understood. Since every compact metrisable space is a quotient of the Cantor set $\Delta:=2^{\mathbb{N}}$, the algebra $C(\Delta)$ of continuous functions on the Cantor set is universal for separable commutative C*-algebras: every such algebra embeds isometrically as a subalgebra of $C(\Delta)$. In contrast, there is no separable C*-algebra that is universal for all separable C*-algebras (commutative or not), as can be deduced from [4, Proposition 2.6].
@@
Our argument is unified and uses only the ordinary projective tensor product. Given a bounded bilinear form $\beta$, we build a separable test algebra $A(\beta)$ whose multiplication remembers $\beta$. Any homomorphic embedding $\phi:A(\beta)\to B$ forces the linearised form $\widetilde{\beta}$ to factor through the fixed separable space $G:=B\widehat{\otimes}_{\!\pi}B$. Choosing $\beta$ so that $\widetilde{\beta}$ (equivalently, a composition $S\circ\widetilde{\beta}_{U}$ defined below) does not factor through $G$, by Johnson–Szankowski [3], yields a contradiction. In the commutative case we pick $\beta$ symmetric so that $A(\beta)$ itself is commutative, but we still linearise via $X\widehat{\otimes}_{\!\pi}X$ and avoid radicals and symmetric tensor products entirely.
We work over $\mathbb{C}$ however the results remain valid for real Banach algebras too. A homomorphism between Banach algebras means a bounded algebra homomorphism. An embedding is an injective homomorphism that is a topological isomorphism onto its closed range.
@@
We present a single proof that covers all four cases. Contractivity plays no rôle beyond boundedness.
@@ 2. Preliminaries
We refer to [6, Chapter 2] for all facts related to the projective tensor product of Banach spaces and to [1, Chapter 2] for Banach-algebraic aspects thereof. For a Banach space $X$ and an integer $n$, the ($n$-fold) projective tensor product $X^{\widehat{\otimes}_{\pi}n}$ (or simply $X\widehat{\otimes}_{\pi}X$ for $n=2$) is defined as the completion of the algebraic tensor product $X^{\otimes n}$ under the projective norm.
@@
The well-definiteness of the operator $u_{1}\otimes_{\pi}u_{2}$ is standard consequence of the universality of the projective tensor product and the norm estimate follows from reasonability of the projective crossnorm. For the identity, evaluate both sides on elementary tensors:
@@
We include the following well-known result for the sake of completeness in order to identify the radical of the test algebras we are about to construct, even though irrelevant for the proof of the main theorem.
@@
Let $A$ be a Banach algebra and let $I\triangleleft A$ be a (two‑sided) ideal with $I^{n}=0$ for some $n\in\mathbb{N}$. Then $I\subseteq\operatorname{rad}(A)$. Moreover, if $A/I$ is semisimple (equivalently $\operatorname{rad}(A/I)=0$), then $\operatorname{rad}(A)\subseteq I$.
@@
If $a\in I$ then $(1-a)(1+a+\dots+a^{n-1})=1$, so 1a1-a is invertible and $a\in\operatorname{rad}(A)$. The quotient statement is standard; see e.g., [1, §3.3]. ∎
@@
This argument uses only that $J$ is nilpotent, so it holds whether or not $\beta$ is symmetric.∎
The key factorisation identity holds without assuming $B$ is commutative.
@@
In $A(\beta)$, $(0,x,0)(0,x^{\prime},0)=(0,0,\beta(x,x^{\prime}))$; applying $\phi$ gives $\mu_{B}(i(x),i(x^{\prime}))=j(\beta(x,x^{\prime}))$. Linearising via Theorem 2.1 and Lemma 2.2 yields the identity. If $\phi$ is an embedding with inverse $\psi$ on its range, then $\|y\|=\|\psi(\phi(0,0,y))\|\leqslant\|\psi\|\,\|j(y)\|$, so $j$ is bounded below. ∎
Finally, we isolate the lifting step in a form that uses a symmetric $\beta_{U}$ but always linearises through $X\widehat{\otimes}_{\!\pi}X$.
@@
Let $G$ be a separable Banach space, and let $U,Y$ be Banach spaces. Suppose $S:U\to Y$ is a compact operator which does not factor through $G$. Set $X:=U\oplus_{1}\mathbb{C}$ and define the symmetric bilinear map
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda u^{\prime}+\lambda^{\prime}u.$
Let $E:=X\widehat{\otimes}_{\!\pi}X$ and let $\widetilde{\beta}_{U}:E\to U$ be the (unsymmetric) linearisation of $\beta_{U}$. Then $\widetilde{\beta}_{U}$ is surjective and admits a bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(u,0)\otimes(0,1)$. Consequently, for $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$ we have:
@@
For $u\in U$, $\widetilde{\beta}_{U}\big((u,0)\otimes(0,1)\big)=\beta_{U}\big((u,0),(0,1)\big)=u.$ Thus $\widetilde{\beta}_{U}$ is onto and $\sigma$ is a right inverse with $\|\sigma(u)\|_{\pi}\leqslant\|(u,0)\|\,\|(0,1)\|=\|u\|$. The two bullet points follow immediately. ∎
3. Proof of the main theorem